OpenStudy (anonymous):

Come up with the specific values of coordinates (x,y) of the highest point on the curve parameterized by [x(t), y(t)] = [te^t, 1-t^2] from t being negative infinty to infinity

OpenStudy (amistre64):

so we should be able to find out whne the slope of the line tangent to the curve is zero

OpenStudy (amistre64):

we can either take the derivative as is; or convert this to an x(y) deal and find when the slope is undefined

OpenStudy (anonymous):

so dy/dx?

OpenStudy (amistre64):

yes, but lets name is dx/dt and dy/dt just to be proper :)

OpenStudy (anonymous):

\[-2t \div t ^{2}e ^{t}\]

OpenStudy (anonymous):

and set that to 0?

OpenStudy (amistre64):

the set to zero part is good, but lets see fif we can work on that derivatives better

OpenStudy (amistre64):

r = < te^t, 1-t^2> the derivative of a componentwise function is the derivative of the components ....

OpenStudy (amistre64):

r' = <t' e^t+t e'^t , (1-t^2)'> r' = <e^t+t e^t , -2t> r' = <e^t(1+t) , -2t> when r' = <0,0> we should be good

OpenStudy (amistre64):

our critical point to chk, is (-1,0) do you see why?

OpenStudy (anonymous):

because at point -1 thats the max x value? because e^0 = 1

OpenStudy (amistre64):

:) remember that we found t values that should suffice; we need to plug these into our original set up to find the actual point that is defined

OpenStudy (amistre64):

that might be in error, but it sounds good lol

OpenStudy (anonymous):

lol im still a bit confused on what to do!!! so do i use dy/dx??

OpenStudy (anonymous):

or just dy/dt and set that equal to 0?

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=y%3D+te%5Et+and+x%3D1-t%5E2 you set dx/dt = 0 and set dy/dt = zero this setup relies upon "t" for its values. x(t) and y(t) are function the depend on "t" to define them

OpenStudy (amistre64):

the point -1,0 on the graph is the low point

OpenStudy (anonymous):

ah and thats how u get -1,0? so what would be the high point?

OpenStudy (amistre64):

its a critical point that we need to look at and determine its condition

OpenStudy (amistre64):

the graph in the link shows that we are at a minimum and there there is no maximum other than infinity

OpenStudy (anonymous):

ok and what do u mean by determine its condition, sorry this is taking me a while to understand

OpenStudy (anonymous):

so no highg point exists?

OpenStudy (amistre64):

when we place a condition upon something, then we want to know when it is doing what we want it to do. when the graph is doing a minimum then it is at a low point, a dip in the graph. when the graph is at a maximum, there will be a high point in the graph. the derivative simply tells us when the graph quits moving up and down; it doesnt tell us the condition of high and low.

OpenStudy (anonymous):

so high point does not exist?

OpenStudy (amistre64):

we need the second derivative if we are to NOT guess at this :)

OpenStudy (anonymous):

ok so then u set that equal to 0? i got -2,0

OpenStudy (anonymous):

xe^x + 2e^x

OpenStudy (amistre64):

r' = <e^t(1+t) , -2t> derive it again for "concavity" r'' = <e^t(1+t)+e^t , -2> r'' = <e^t(2+t) , -2> ; this is never at (0,0)

OpenStudy (anonymous):

2,-2?

OpenStudy (anonymous):

err -2,-2 sorry i meant

OpenStudy (anonymous):

asdf;laskjfaf lol so lost

OpenStudy (amistre64):

when t=-2, there might be concavity; its hard for me to say cause I cant be to certain :) but, the value for r' seem to be values of t for each component

OpenStudy (amistre64):

when t=-1, x is at a critical point; when t=0 then y is at a critical point

OpenStudy (amistre64):

when t=-2, then x should be at a turning point in concavity

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

so would be the highest point, like the actual answer?

OpenStudy (turingtest):

amistre are you sure we want r'=<0,0> ? my understanding is a bit different....

OpenStudy (amistre64):

i was until i started trying to fit it to the graph :) but now i have to wonder ....

OpenStudy (turingtest):

I think we need\[\frac{dy}{dx}={{dy\over dt}\over{dx\over dt}}=0\]

OpenStudy (turingtest):

so that's\[\frac{-2t}{e^t(1+t)}=0\implies t=0\]as our only critical point

OpenStudy (anonymous):

ah so would it be 0,1 then if u plug the value of t back into its original functions?

OpenStudy (amistre64):

if got my x andy backwards in me graph :)

OpenStudy (anonymous):

and that would be my high point? 0,1?

OpenStudy (turingtest):

yes, I believe the point in question is <0,1> whether we can show it is a max and not a min I'm not sure about

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=x%3D+te%5Et+and+y%3D1-t%5E2%2C+t+from+-2+to+2 i had this thing laying on its side lol

OpenStudy (anonymous):

well yeah u can u choose a point before and a point after and see that it is less then 0,1

OpenStudy (turingtest):

yeah, but I wanted to do it with concavity oh well, it is clearly the max, so there ya go :D

OpenStudy (turingtest):

yeah, second derivative is\[\frac{-2}{e^t(1+t)}<0\]so there we can show it with concavity

OpenStudy (anonymous):

thank u guys!

OpenStudy (amistre64):

good job :)